# A Dynamical Theory of Overfitting and Generalization in Large Two-Layer Networks

### v2

@@ Appendix I Lower bounding the overfitting timescale
$\frac{{\rm d}a_{i}(t)}{{\rm d}t}=$ $\;\frac{\overline{\alpha}}{m}{\mathbb{E}}\big{\{}E(t)\,\sigma(r_{i}(t))\big{\}}\,$, (J.1)
$\nu_{i}(t)=$ $\;\frac{\overline{\alpha}}{m}a_{i}(t){\mathbb{E}}\big{\{}E(t)\sigma^{\prime}(r_{i}(t))r_{i}(t)\big{\}}\,$, (J.2)
$\frac{{\rm d}w^{e}_{i}(t)}{{\rm d}t}=$ $\;-\nu_{i}(t)w^{e}_{i}(t)-\frac{1}{m}\sum_{l=1}^{m}\int_{0}^{t}M_{i,l}(t,s)w^{e}_{l}(s)\,{\rm d}s$ (J.3)
@@
$r_{i}(t)=$ $\;\frac{1}{m}\sum_{l=1}^{m}\int_{0}^{t}R_{il}(t,s)\,a_{l}(s)E(s)\,\sigma^{\prime}(r_{l}(s))\,{\rm d}s+\xi_{i}(t)\,$, ${\boldsymbol{\xi}}\sim{\sf GP}(0,{\boldsymbol{C}})\,$, (J.4)
$E(t):=$ $\;y-\frac{1}{m}\sum_{l=1}^{m}a_{l}(t)\sigma(r_{l}(t))\,.$ (J.5)
@@
$M_{ij}(t,s)=$ $\overline{\alpha}{\mathbb{E}}\{S_{ij}(t)\}\,\delta(t-s)+\overline{\alpha}\sum_{l=1}^{m}{\mathbb{E}}\Big{\{}S_{il}(t)\frac{\partial r_{l}(t)}{\partial\xi_{j}(s)}\big{\}}\,$, (J.6)
$M_{ij}(t,*)=$ $-\overline{\alpha}\frac{a_{i}(t)}{m}{\mathbb{E}}\big{\{}\sigma^{\prime}(r_{i}(t))\nabla_{j}\varphi({\boldsymbol{r}}_{0})\big{\}}+\frac{\overline{\alpha}}{m}\sum_{l=1}^{m}{\mathbb{E}}\Big{\{}S_{il}(t)\frac{\partial r_{l}(t)}{\partial r_{0,j}}\big{\}}\,$, (J.7)
$C^{E}_{i,j}(t,s)$ $=\overline{\alpha}\frac{a_{i}(t)a_{j}(s)}{m^{2}}{\mathbb{E}}\left\{E(t)E(s)\,\sigma^{\prime}(r_{i}(t))\sigma^{\prime}(r_{j}(s))\right\}\,$, (J.8)
$S_{ij}(t):=$ $-a_{i}(t)\,E(t)\sigma^{\prime\prime}(r_{i}(t))\delta_{ij}+\frac{a_{i}(t)a_{j}(t)}{m}\sigma^{\prime}(r_{i}(t))\sigma^{\prime}(r_{j}(t))\,$, (J.9)
@@
$C_{ij}(t,s)$ $={\mathbb{E}}\Big{\{}w^{e}_{i}(t)w^{e}_{j}(s)\Big{\}}\,$, (J.10)
$R_{ij}(t,s)$ $={\mathbb{E}}\Big{\{}\frac{\partial w^{e}_{i}(t)}{\partial\eta_{j}(s)}\Big{\}}\,.$ (J.11)
@@
$\frac{{\rm d}\phantom{t}}{{\rm d}t}\frac{\partial w^{e}_{i}(t)}{\partial\eta_{j}(s)}=$ $\;-\nu_{i}(t)\frac{\partial w^{e}_{i}(t)}{\partial\eta_{j}(s)}-\frac{1}{m}\sum_{l=1}^{m}\int_{s}^{t}M_{i,l}(t,t^{\prime})\frac{\partial w^{e}_{l}(t^{\prime})}{\partial\eta_{j}(s)}\,{\rm d}t\,$, (J.12)
$\frac{\partial r_{i}(t)}{\partial\xi_{j}(s)}=$ $-\frac{1}{m}\sum_{l,q=1}^{m}\int_{s}^{t}R_{il}(t,t^{\prime})S_{lq}(t^{\prime})\Big{[}\frac{\partial r_{q}(t^{\prime})}{\partial\xi_{j}(s)}+\delta_{qj}\delta(t^{\prime}-s)\Big{]}\,{\rm d}t^{\prime}\,$, (J.13)
$\frac{\partial r_{i}(t)}{\partial r_{0,j}}=$ $\frac{1}{m}\sum_{l=1}^{m}\int_{0}^{t}R_{il}(t,s)a_{l}(s)\sigma^{\prime}(r_{l}(s))\nabla_{j}\varphi({\boldsymbol{r}}_{0})\,{\rm d}s-\frac{1}{m}\sum_{l,q=1}^{m}\int_{0}^{t}R_{il}(t,s)S_{lq}(s)\frac{\partial r_{q}(s)}{\partial r_{0,j}}{\rm d}s\,$, (J.14)
@@
Appendix K Derivation of the dynamical mean field theory equations
@@
$\begin{split}A=i\sum_{l=1}^{m}\int\tilde{a}_{l}(t)\left[d\frac{{\rm d}a_{l}(t)}{{\rm d}t}+n\frac{\partial\widehat{\mathscrsfs{R}}_{n}}{\partial a_{l}(t)}\right]{\rm d}t+i\sum_{l=1}^{m}\int\langle{\hat{\boldsymbol{w}}}_{l}(t),d\frac{{{\boldsymbol{w}}}_{l}(t)}{{\rm d}t}+d\nu_{i}(t){\boldsymbol{w}}_{i}(t)+n\frac{\partial{\widehat{\mathscrsfs{R}}_{n}}}{\partial{\boldsymbol{w}}_{l}(t)}\rangle{\rm d}t\>.\end{split}$ (K.2)
Eq. (K.2) can be rewritten by introducing Grassmann variables [ZJ21]. Call $\hat{a}=(t_{a},\theta_{a})$ a supertime coordinate, with $\theta_{a}$ a Grassmann variable. Define, with a slight abuse of notation
$\begin{split}{\boldsymbol{w}}_{l}(\hat{a})&={\boldsymbol{w}}_{l}(t_{a})+i\theta_{a}{\hat{\boldsymbol{w}}}_{l}\\ a_{l}(\hat{a})&=a_{l}(t_{a})+i\theta_{a}\tilde{a}_{l}(t_{a})\ \ \ \ l\leq m\>.\end{split}$ (K.3)
Eq. (K.2) can be written as
@@
$Z_{dyn}=\int D(\underline{a},\tilde{Q},R)\exp\left[S_{dyn}(\underline{a},\tilde{Q},R)\right]\>.$ (K.6)
@@
In the large $d$ limit, fixing $m$ and $\overline{\alpha}$, the path integral in Eq. (K.5) concentrates on its saddle point. The corresponding equations are
@@
Eqs. (K.9)-(K.12) contain all the information about the dynamics. In order to fully specify the behavior of physical quantities such has the train and test error, it is useful to unfold the Grassmann structure of Eqs. (K.9)-(K.12).
K.1 Unfolding the Grassmann structure
@@
$\begin{split}r_{\alpha}(\hat{a})&=r_{\alpha}(t_{a})\\ a_{\alpha}(\hat{a})&=a_{\alpha}(t_{a})\\ Q_{\alpha\beta}(\hat{a},\hat{b})&=C_{\alpha\beta}(t_{a},t_{b})+\theta_{a}R_{\beta\alpha}(t_{b},t_{a})+\theta_{b}R_{\alpha\beta}(t_{a},t_{b})\\ ({\boldsymbol{I}}+\Sigma)^{-1}(\hat{a},\hat{b})&=C_{A}(t_{a},t_{b})+\theta_{b}R_{A}(t_{a},t_{b})+\theta_{a}R_{A}(t_{b},t_{a})\>.\end{split}$ (K.13)
Plugging this parametrization into the saddle point equations we get that the correlators in Eqs. (K.13) satisfy the following DMFT equations
$\begin{split}\frac{{\rm d}a_{\alpha}(t)}{{\rm d}t}&=-\frac{\overline{\alpha}}{m}\int_{0}^{t}R_{A}(t,s)\left[\frac{1}{m}\sum_{l=1}^{m}a_{l}(s)h\left[C_{l\alpha}(s,t)\right]-\hat{\varphi}(r_{\alpha}(t))\right]{\rm d}s\\ &-\frac{\overline{\alpha}}{m}\int_{0}^{t}C_{A}(t,s)\frac{1}{m}\sum_{l=1}^{m}a_{l}(s)h^{\prime}[C_{l\alpha}(s,t)]R_{\alpha l}(t,s){\rm d}s\end{split}$ (K.14)
$\begin{split}\frac{{\rm d}r_{\alpha}(t)}{{\rm d}t}&=-\nu_{\alpha}(t)r_{\alpha}(t)+\frac{\overline{\alpha}}{m}a_{\alpha}(t)\hat{\varphi}^{\prime}(r_{\alpha}(t))\int_{0}^{t}R_{A}(t,s){\rm d}s\\ &-\frac{a_{\alpha}(t)}{m}\sum_{\gamma=1}^{m}\int_{0}^{t}M^{R}_{\alpha\gamma}(t,s)a_{\gamma}(s)r_{\gamma}(s){\rm d}s\end{split}$ (K.15)
$\begin{split}\frac{\partial C_{\alpha\beta}(t_{a},t_{b})}{\partial t_{a}}&=-\nu_{\alpha}(t_{a})C_{\alpha\beta}(t_{a},t_{b})+\frac{\overline{\alpha}}{m}a_{\alpha}(t_{a})\hat{\varphi}^{\prime}(r_{\alpha}(t_{a}))r_{\beta}(t_{b})\int_{0}^{t_{a}}R_{A}(t_{a},s){\rm d}s\\ &-\frac{a_{\alpha}(t_{a})}{m}\sum_{\gamma=1}^{m}\int_{0}^{t_{a}}M^{R}_{\alpha\gamma}(t_{a},s)a_{\gamma}(s)C_{\gamma\beta}(s,t_{b}){\rm d}s\\ &-\frac{a_{\alpha}(t_{a})}{m}\sum_{\gamma=1}^{m}\int_{0}^{t_{b}}M^{C}_{\alpha\gamma}(t_{a},s)a_{\gamma}(s)R_{\beta\gamma}(t_{b},s){\rm d}s\end{split}$ (K.16)
$\begin{split}\frac{\partial R_{\alpha\beta}(t_{a},t_{b})}{\partial t_{a}}&=-\nu_{\alpha}(t_{a})R_{\alpha\beta}(t_{a},t_{b})+\delta_{\alpha\beta}(t_{a}-t_{b})\\ &-\frac{a_{\alpha}(t_{a})}{m}\sum_{\gamma=1}^{m}\int_{t_{b}}^{t_{a}}M^{R}_{\alpha\gamma}(t_{a},s)a_{\gamma}(s)R_{\gamma\beta}(s,t_{b}){\rm d}s\>.\end{split}$ (K.17)
@@
$\begin{split}M^{R}_{\alpha\gamma}(t,s)&=\frac{\overline{\alpha}}{m}\left[R_{A}(t,s)h^{\prime}(C_{\alpha\gamma}(t,s))+C_{A}(t,s)h^{\prime\prime}(C_{\alpha\gamma}(t,s))R_{\alpha\gamma}(t,s)\right]\\ M^{C}_{\alpha\gamma}(t,s)&=\frac{\overline{\alpha}}{m}C_{A}(t,s)h^{\prime}(C_{\alpha\gamma}(t,s))\>.\end{split}$ (K.18)
The kernels in Eq. (K.18) depend on $R_{A}$ and $C_{A}$ that are defined in Eqs. (K.13). The corresponding equations are
$\begin{split}&\int_{t^{\prime}}^{t}\left[\delta(t-s)+\Sigma_{R}(t,s)\right]R_{A}(s,t^{\prime}){\rm d}s=\delta(t-t^{\prime})\\ &\int_{0}^{t}\left[\delta(t-s)+\Sigma_{R}(t,s)\right]C_{A}(s,t^{\prime}){\rm d}s+\int_{0}^{t^{\prime}}\Sigma_{C}(t,s)R_{A}(t^{\prime},s){\rm d}s=0\end{split}$ (K.19)
@@
$\begin{split}\Sigma_{C}(t,s)&=\tau^{2}+h_{t}(1)+\frac{1}{m^{2}}\sum_{ll^{\prime}=1}a_{l}(t)a_{l^{\prime}}(s)h[C_{ll^{\prime}}(t,s)]\\ &-\frac{1}{m}\sum_{l=1}^{m}a_{l}(t)\hat{\varphi}(r_{l}(t))-\frac{1}{m}\sum_{l=1}^{m}a_{l}(s)\hat{\varphi}(r_{l}(s))\\ \Sigma_{R}(t,s)&=\frac{1}{m^{2}}\sum_{ll^{\prime}=1}^{m}a_{l}(t)a_{l^{\prime}}(s)h^{\prime}[C_{ll^{\prime}}(t,s)]R_{ll^{\prime}}(t,s)\>.\end{split}$ (K.20)
@@
$\begin{split}\nu_{\alpha}(t_{a})&=\frac{\overline{\alpha}}{km}\sum_{\tau=1}^{k}a_{\alpha}(t_{a})\hat{\varphi}^{\prime}(r_{\tau\alpha}(t_{a}))r_{\tau\alpha}(t_{a})\int_{0}^{t_{a}}R_{A}(t_{a},s){\rm d}s\\ &-\frac{a_{\alpha}(t_{a})}{m}\sum_{\gamma=1}^{m}\int_{0}^{t_{a}}M^{R}_{\alpha\gamma}(t_{a},s)a_{\gamma}(s)C_{\gamma\alpha}(s,t_{a}){\rm d}s\\ &-\frac{a_{\alpha}(t_{a})}{m}\sum_{\gamma=1}^{m}\int_{0}^{t_{a}}M^{C}_{\alpha\gamma}(t_{a},s)a_{\gamma}(s)R_{\gamma\alpha}(t_{a},s){\rm d}s\end{split}$ (K.21)
@@
$\begin{split}\frac{{\rm d}C_{\alpha\beta}(t_{a},t_{a})}{{\rm d}t_{a}}&=\lim_{t^{\prime}\to t_{a}}\left[\frac{\partial C_{\alpha\beta}(t_{a},t^{\prime})}{\partial t_{a}}+\frac{\partial C_{\beta\alpha}(t_{a},t^{\prime})}{\partial t_{a}}\right]\>.\end{split}$ (K.22)
@@
$\begin{split}r_{\alpha}(0)&=0\hskip 28.45274pt\forall\alpha=1,\ldots,m\\ C_{\alpha\neq\beta}(0,0)&=0\hskip 28.45274pt\forall\alpha\neq\beta=1,\ldots,m\\ C_{\alpha\alpha}(0,0)&=1\hskip 28.45274pt\forall\alpha=1,\ldots,m\\ R_{\alpha\beta}(0,0)&=0\hskip 28.45274pt\forall\alpha,\beta=1,\ldots,m\>.\end{split}$ (K.23)
@@
$e_{\mbox{{\rm\footnotesize tr}}}(t)=\lim_{d\to\infty}\widehat{\mathscrsfs{R}}_{n}(t)\>.$ (K.24)
A simple way to derive the expression of $e_{\mbox{{\rm\footnotesize tr}}}$ as a function of the solution of the DMFT equations in the $d\to\infty$ limit is to consider a deformation of Eq. (K.5) which consists in replacing
@@
$e_{\mbox{{\rm\footnotesize tr}}}(t)=-\int{\rm d}\theta_{a}\left.\frac{\delta}{\delta P(\hat{a})}\ln Z_{dyn}[P]\right|_{P=1}\>.$ (K.26)
Note that the deformed dynamical partition function $Z_{dyn}[P]$ does not equal 1 for generic $P$ so that the formula above makes perfectly sense. The deformation of the partition function produces a deformation of $S_{dyn}$ in Eq. (K.7) which consist in replacing
@@
$e_{\mbox{{\rm\footnotesize tr}}}(t)=\frac{1}{2}\int_{0}^{t}\left[R_{A}(t,s)\Sigma_{C}(t,s)+C_{A}(t,s)\Sigma_{R}(t,s)\right]{\rm d}s\>.$ (K.28)
@@
$\begin{split}e_{\mbox{\rm\footnotesize ts}}(t)&=\lim_{d\to\infty}\frac{1}{2}{\mathbb{E}}\left[\left(y_{\rm new})-y^{(s)}_{\rm new}\right)^{2}\right]\\ &=\frac{1}{2}\left[\tau^{2}+\frac{1}{k}h_{t}(1)+\frac{1}{m^{2}}\sum_{ll^{\prime}}^{m}h[C_{ll^{\prime}}(t,t)]-2\frac{1}{m}\sum_{l}^{m}\hat{\varphi}(r_{l}(t))\right]\>.\end{split}$ (K.29)
The average in Eq. (K.29) is performed over the training set and an additional datapoint, not presented in the training set and having the same statistical structure.

### v3

@@ Appendix I Details about SGD simulations
$\frac{{\rm d}a_{i}(t)}{{\rm d}t}=$ $\;\frac{\overline{\alpha}}{m}{\mathbb{E}}\big\{E(t)\,\sigma(r_{i}(t))\big\}\,$, (K.1)
$\nu_{i}(t)=$ $\;\frac{\overline{\alpha}}{m}a_{i}(t){\mathbb{E}}\big\{E(t)\sigma^{\prime}(r_{i}(t))r_{i}(t)\big\}\,$, (K.2)
$\frac{{\rm d}w^{e}_{i}(t)}{{\rm d}t}=$ $\;-\nu_{i}(t)w^{e}_{i}(t)-\frac{1}{m}\sum_{l=1}^{m}\int_{0}^{t}M_{i,l}(t,s)w^{e}_{l}(s)\,{\rm d}s$ (K.3)
@@
$r_{i}(t)=$ $\;\frac{1}{m}\sum_{l=1}^{m}\int_{0}^{t}R_{il}(t,s)\,a_{l}(s)E(s)\,\sigma^{\prime}(r_{l}(s))\,{\rm d}s+\xi_{i}(t)\,$, ${\boldsymbol{\xi}}\sim{\sf GP}(0,{\boldsymbol{C}})\,$, (K.4)
$E(t):=$ $\;y-\frac{1}{m}\sum_{l=1}^{m}a_{l}(t)\sigma(r_{l}(t))\,.$ (K.5)
@@
$M_{ij}(t,s)=$ $\overline{\alpha}{\mathbb{E}}\{S_{ij}(t)\}\,\delta(t-s)+\overline{\alpha}\sum_{l=1}^{m}{\mathbb{E}}\Big\{S_{il}(t)\frac{\partial r_{l}(t)}{\partial\xi_{j}(s)}\big\}\,$, (K.6)
$M_{ij}(t,*)=$ $-\overline{\alpha}\frac{a_{i}(t)}{m}{\mathbb{E}}\big\{\sigma^{\prime}(r_{i}(t))\nabla_{j}\varphi({\boldsymbol{r}}_{0})\big\}+\frac{\overline{\alpha}}{m}\sum_{l=1}^{m}{\mathbb{E}}\Big\{S_{il}(t)\frac{\partial r_{l}(t)}{\partial r_{0,j}}\big\}\,$, (K.7)
$C^{E}_{i,j}(t,s)$ $=\overline{\alpha}\frac{a_{i}(t)a_{j}(s)}{m^{2}}{\mathbb{E}}\left\{E(t)E(s)\,\sigma^{\prime}(r_{i}(t))\sigma^{\prime}(r_{j}(s))\right\}\,$, (K.8)
$S_{ij}(t):=$ $-a_{i}(t)\,E(t)\sigma^{\prime\prime}(r_{i}(t))\delta_{ij}+\frac{a_{i}(t)a_{j}(t)}{m}\sigma^{\prime}(r_{i}(t))\sigma^{\prime}(r_{j}(t))\,$, (K.9)
@@
$C_{ij}(t,s)$ $={\mathbb{E}}\Big\{w^{e}_{i}(t)w^{e}_{j}(s)\Big\}\,$, (K.10)
$R_{ij}(t,s)$ $={\mathbb{E}}\Big\{\frac{\partial w^{e}_{i}(t)}{\partial\eta_{j}(s)}\Big\}\,.$ (K.11)
@@
$\frac{{\rm d}\phantom{t}}{{\rm d}t}\frac{\partial w^{e}_{i}(t)}{\partial\eta_{j}(s)}=$ $\;-\nu_{i}(t)\frac{\partial w^{e}_{i}(t)}{\partial\eta_{j}(s)}-\frac{1}{m}\sum_{l=1}^{m}\int_{s}^{t}M_{i,l}(t,t^{\prime})\frac{\partial w^{e}_{l}(t^{\prime})}{\partial\eta_{j}(s)}\,{\rm d}t\,$, (K.12)
$\frac{\partial r_{i}(t)}{\partial\xi_{j}(s)}=$ $-\frac{1}{m}\sum_{l,q=1}^{m}\int_{s}^{t}R_{il}(t,t^{\prime})S_{lq}(t^{\prime})\Big[\frac{\partial r_{q}(t^{\prime})}{\partial\xi_{j}(s)}+\delta_{qj}\delta(t^{\prime}-s)\Big]\,{\rm d}t^{\prime}\,$, (K.13)
$\frac{\partial r_{i}(t)}{\partial r_{0,j}}=$ $\frac{1}{m}\sum_{l=1}^{m}\int_{0}^{t}R_{il}(t,s)a_{l}(s)\sigma^{\prime}(r_{l}(s))\nabla_{j}\varphi({\boldsymbol{r}}_{0})\,{\rm d}s-\frac{1}{m}\sum_{l,q=1}^{m}\int_{0}^{t}R_{il}(t,s)S_{lq}(s)\frac{\partial r_{q}(s)}{\partial r_{0,j}}{\rm d}s\,$, (K.14)
@@
Appendix L Derivation of the dynamical mean field theory equations
@@
$\begin{split}A=i\sum_{l=1}^{m}\int\tilde{a}_{l}(t)\left[d\frac{{\rm d}a_{l}(t)}{{\rm d}t}+n\frac{\partial\widehat{\mathscrsfs{R}}_{n}}{\partial a_{l}(t)}\right]{\rm d}t+i\sum_{l=1}^{m}\int\langle{\hat{\boldsymbol{w}}}_{l}(t),d\frac{{{\boldsymbol{w}}}_{l}(t)}{{\rm d}t}+d\nu_{i}(t){\boldsymbol{w}}_{i}(t)+n\frac{\partial{\widehat{\mathscrsfs{R}}_{n}}}{\partial{\boldsymbol{w}}_{l}(t)}\rangle{\rm d}t\>.\end{split}$ (L.2)
Eq. (L.2) can be rewritten by introducing Grassmann variables [ZJ21]. Call $\hat{a}=(t_{a},\theta_{a})$ a supertime coordinate, with $\theta_{a}$ a Grassmann variable. Define, with a slight abuse of notation
$\begin{split}{\boldsymbol{w}}_{l}(\hat{a})&={\boldsymbol{w}}_{l}(t_{a})+i\theta_{a}{\hat{\boldsymbol{w}}}_{l}\\ a_{l}(\hat{a})&=a_{l}(t_{a})+i\theta_{a}\tilde{a}_{l}(t_{a})\ \ \ \ l\leq m\>.\end{split}$ (L.3)
Eq. (L.2) can be written as
@@
$Z_{dyn}=\int D(\underline{a},\tilde{Q},R)\exp\left[S_{dyn}(\underline{a},\tilde{Q},R)\right]\>.$ (L.6)
@@
In the large $d$ limit, fixing $m$ and $\overline{\alpha}$, the path integral in Eq. (L.5) concentrates on its saddle point. The corresponding equations are
@@
Eqs. (L.9)-(L.12) contain all the information about the dynamics. In order to fully specify the behavior of physical quantities such has the train and test error, it is useful to unfold the Grassmann structure of Eqs. (L.9)-(L.12).
L.1 Unfolding the Grassmann structure
@@
$\begin{split}r_{\alpha}(\hat{a})&=r_{\alpha}(t_{a})\\ a_{\alpha}(\hat{a})&=a_{\alpha}(t_{a})\\ Q_{\alpha\beta}(\hat{a},\hat{b})&=C_{\alpha\beta}(t_{a},t_{b})+\theta_{a}R_{\beta\alpha}(t_{b},t_{a})+\theta_{b}R_{\alpha\beta}(t_{a},t_{b})\\ ({\boldsymbol{I}}+\Sigma)^{-1}(\hat{a},\hat{b})&=C_{A}(t_{a},t_{b})+\theta_{b}R_{A}(t_{a},t_{b})+\theta_{a}R_{A}(t_{b},t_{a})\>.\end{split}$ (L.13)
Plugging this parametrization into the saddle point equations we get that the correlators in Eqs. (L.13) satisfy the following DMFT equations
$\begin{split}\frac{{\rm d}a_{\alpha}(t)}{{\rm d}t}&=-\frac{\overline{\alpha}}{m}\int_{0}^{t}R_{A}(t,s)\left[\frac{1}{m}\sum_{l=1}^{m}a_{l}(s)h\left[C_{l\alpha}(s,t)\right]-\hat{\varphi}(r_{\alpha}(t))\right]{\rm d}s\\ &-\frac{\overline{\alpha}}{m}\int_{0}^{t}C_{A}(t,s)\frac{1}{m}\sum_{l=1}^{m}a_{l}(s)h^{\prime}[C_{l\alpha}(s,t)]R_{\alpha l}(t,s){\rm d}s\end{split}$ (L.14)
$\begin{split}\frac{{\rm d}r_{\alpha}(t)}{{\rm d}t}&=-\nu_{\alpha}(t)r_{\alpha}(t)+\frac{\overline{\alpha}}{m}a_{\alpha}(t)\hat{\varphi}^{\prime}(r_{\alpha}(t))\int_{0}^{t}R_{A}(t,s){\rm d}s\\ &-\frac{a_{\alpha}(t)}{m}\sum_{\gamma=1}^{m}\int_{0}^{t}M^{R}_{\alpha\gamma}(t,s)a_{\gamma}(s)r_{\gamma}(s){\rm d}s\end{split}$ (L.15)
$\begin{split}\frac{\partial C_{\alpha\beta}(t_{a},t_{b})}{\partial t_{a}}&=-\nu_{\alpha}(t_{a})C_{\alpha\beta}(t_{a},t_{b})+\frac{\overline{\alpha}}{m}a_{\alpha}(t_{a})\hat{\varphi}^{\prime}(r_{\alpha}(t_{a}))r_{\beta}(t_{b})\int_{0}^{t_{a}}R_{A}(t_{a},s){\rm d}s\\ &-\frac{a_{\alpha}(t_{a})}{m}\sum_{\gamma=1}^{m}\int_{0}^{t_{a}}M^{R}_{\alpha\gamma}(t_{a},s)a_{\gamma}(s)C_{\gamma\beta}(s,t_{b}){\rm d}s\\ &-\frac{a_{\alpha}(t_{a})}{m}\sum_{\gamma=1}^{m}\int_{0}^{t_{b}}M^{C}_{\alpha\gamma}(t_{a},s)a_{\gamma}(s)R_{\beta\gamma}(t_{b},s){\rm d}s\end{split}$ (L.16)
$\begin{split}\frac{\partial R_{\alpha\beta}(t_{a},t_{b})}{\partial t_{a}}&=-\nu_{\alpha}(t_{a})R_{\alpha\beta}(t_{a},t_{b})+\delta_{\alpha\beta}(t_{a}-t_{b})\\ &-\frac{a_{\alpha}(t_{a})}{m}\sum_{\gamma=1}^{m}\int_{t_{b}}^{t_{a}}M^{R}_{\alpha\gamma}(t_{a},s)a_{\gamma}(s)R_{\gamma\beta}(s,t_{b}){\rm d}s\>.\end{split}$ (L.17)
@@
$\begin{split}M^{R}_{\alpha\gamma}(t,s)&=\frac{\overline{\alpha}}{m}\left[R_{A}(t,s)h^{\prime}(C_{\alpha\gamma}(t,s))+C_{A}(t,s)h^{\prime\prime}(C_{\alpha\gamma}(t,s))R_{\alpha\gamma}(t,s)\right]\\ M^{C}_{\alpha\gamma}(t,s)&=\frac{\overline{\alpha}}{m}C_{A}(t,s)h^{\prime}(C_{\alpha\gamma}(t,s))\>.\end{split}$ (L.18)
The kernels in Eq. (L.18) depend on $R_{A}$ and $C_{A}$ that are defined in Eqs. (L.13). The corresponding equations are
$\begin{split}&\int_{t^{\prime}}^{t}\left[\delta(t-s)+\Sigma_{R}(t,s)\right]R_{A}(s,t^{\prime}){\rm d}s=\delta(t-t^{\prime})\\ &\int_{0}^{t}\left[\delta(t-s)+\Sigma_{R}(t,s)\right]C_{A}(s,t^{\prime}){\rm d}s+\int_{0}^{t^{\prime}}\Sigma_{C}(t,s)R_{A}(t^{\prime},s){\rm d}s=0\end{split}$ (L.19)
@@
$\begin{split}\Sigma_{C}(t,s)&=\tau^{2}+h_{t}(1)+\frac{1}{m^{2}}\sum_{ll^{\prime}=1}a_{l}(t)a_{l^{\prime}}(s)h[C_{ll^{\prime}}(t,s)]\\ &-\frac{1}{m}\sum_{l=1}^{m}a_{l}(t)\hat{\varphi}(r_{l}(t))-\frac{1}{m}\sum_{l=1}^{m}a_{l}(s)\hat{\varphi}(r_{l}(s))\\ \Sigma_{R}(t,s)&=\frac{1}{m^{2}}\sum_{ll^{\prime}=1}^{m}a_{l}(t)a_{l^{\prime}}(s)h^{\prime}[C_{ll^{\prime}}(t,s)]R_{ll^{\prime}}(t,s)\>.\end{split}$ (L.20)
@@
$\begin{split}\nu_{\alpha}(t_{a})&=\frac{\overline{\alpha}}{km}\sum_{\tau=1}^{k}a_{\alpha}(t_{a})\hat{\varphi}^{\prime}(r_{\tau\alpha}(t_{a}))r_{\tau\alpha}(t_{a})\int_{0}^{t_{a}}R_{A}(t_{a},s){\rm d}s\\ &-\frac{a_{\alpha}(t_{a})}{m}\sum_{\gamma=1}^{m}\int_{0}^{t_{a}}M^{R}_{\alpha\gamma}(t_{a},s)a_{\gamma}(s)C_{\gamma\alpha}(s,t_{a}){\rm d}s\\ &-\frac{a_{\alpha}(t_{a})}{m}\sum_{\gamma=1}^{m}\int_{0}^{t_{a}}M^{C}_{\alpha\gamma}(t_{a},s)a_{\gamma}(s)R_{\gamma\alpha}(t_{a},s){\rm d}s\end{split}$ (L.21)
@@
$\begin{split}\frac{{\rm d}C_{\alpha\beta}(t_{a},t_{a})}{{\rm d}t_{a}}&=\lim_{t^{\prime}\to t_{a}}\left[\frac{\partial C_{\alpha\beta}(t_{a},t^{\prime})}{\partial t_{a}}+\frac{\partial C_{\beta\alpha}(t_{a},t^{\prime})}{\partial t_{a}}\right]\>.\end{split}$ (L.22)
@@
$\begin{split}r_{\alpha}(0)&=0\hskip 28.45274pt\forall\alpha=1,\ldots,m\\ C_{\alpha\neq\beta}(0,0)&=0\hskip 28.45274pt\forall\alpha\neq\beta=1,\ldots,m\\ C_{\alpha\alpha}(0,0)&=1\hskip 28.45274pt\forall\alpha=1,\ldots,m\\ R_{\alpha\beta}(0,0)&=0\hskip 28.45274pt\forall\alpha,\beta=1,\ldots,m\>.\end{split}$ (L.23)
@@
$e_{\mbox{{\rm\footnotesize tr}}}(t)=\lim_{d\to\infty}\widehat{\mathscrsfs{R}}_{n}(t)\>.$ (L.24)
A simple way to derive the expression of $e_{\mbox{{\rm\footnotesize tr}}}$ as a function of the solution of the DMFT equations in the $d\to\infty$ limit is to consider a deformation of Eq. (L.5) which consists in replacing
@@
$e_{\mbox{{\rm\footnotesize tr}}}(t)=-\int{\rm d}\theta_{a}\left.\frac{\delta}{\delta P(\hat{a})}\ln Z_{dyn}[P]\right|_{P=1}\>.$ (L.26)
Note that the deformed dynamical partition function $Z_{dyn}[P]$ does not equal 1 for generic $P$ so that the formula above makes perfectly sense. The deformation of the partition function produces a deformation of $S_{dyn}$ in Eq. (L.7) which consist in replacing
@@
$e_{\mbox{{\rm\footnotesize tr}}}(t)=\frac{1}{2}\int_{0}^{t}\left[R_{A}(t,s)\Sigma_{C}(t,s)+C_{A}(t,s)\Sigma_{R}(t,s)\right]{\rm d}s\>.$ (L.28)
@@
$\begin{split}e_{\mbox{\rm\footnotesize ts}}(t)&=\lim_{d\to\infty}\frac{1}{2}{\mathbb{E}}\left[\left(y_{\rm new})-y^{(s)}_{\rm new}\right)^{2}\right]\\ &=\frac{1}{2}\left[\tau^{2}+\frac{1}{k}h_{t}(1)+\frac{1}{m^{2}}\sum_{ll^{\prime}}^{m}h[C_{ll^{\prime}}(t,t)]-2\frac{1}{m}\sum_{l}^{m}\hat{\varphi}(r_{l}(t))\right]\>.\end{split}$ (L.29)
The average in Eq. (L.29) is performed over the training set and an additional datapoint, not presented in the training set and having the same statistical structure.